OpenStudy (anonymous):

Which of the following equations has only one solution? x^2 + 4x + 4 = 0 x^2 + x = 0 x^2 - 1 = 0

OpenStudy (vishweshshrimali5):

Best would be that you find out solutions for each

OpenStudy (vishweshshrimali5):

Lets start from the first one

OpenStudy (vishweshshrimali5):

\[\large{x^2 + 4x + 4 = x^2 + 2\times 2 \times x + 2^2}\] Compare this to: \[\large{(a+b)^2 = a^2 + 2\times a\times b + b^2}\]